Back to the NIKA2 main page

Daily Reports

Feb 8th, Tuesday

15:45-21:00; Angel: DAQ started and tuning with 48/7/54 bad KIDs for ar1/ar2/ar3. No evidence of correlated spikes in the noise spectra for all arrays. Tau(225 GHz) between 0.20 and 0.25 throughout the shift with some spikes. Pointing on Uranus (1.5, -11.6), larger in elevation than expected. NIKA2 mirrors deployment re-checked. Best focus -1.0 mm. Pointing corrections are stable at this position. Beam sizes: 12.9"/18" because of a slight ellipticity; beam shapes were quite reasonable. Good atmosphere stability: ~2 Jy/b at 2 mm. Checking pointing corrections (6 values) in ~70 < Az. < 120 deg, and ~45 < El. < 70 deg. We get corrections <Az.> = 3.2 +/- 1.4 and <El.> = -8.9 +/- 3.0. Refocusing and checking stability (~1-2 Jy/b at 2 mm). Science observations started with project 110-21 (2 reps. on NGC1333-C7). Then, 2 reps. on B213E (GASTON). Even more stable after sunset. The shift was finished with the observation of target WB89-793 of project 137-21. Stuck scans: 20220208s122 and s159.

21:00-03:00; Longji: We start the shift with a series of pointing, focus, and calibration scans on Uranus and a skydip [beam map on Uranus: 20220208s191; calib1scan: s192; skydip: s193]. After that we made 2 repetitions on NGC3938 for IMEGIN (160-16). We started observation on COSMOS for N2CLS (192-16) at midnight. The opacity from the tau meter was slightly spiky at the beginning of the observation but became stable after 23:00, with values of ~0.23. The beam and focus were also stable throughout the shift.

Feb 9th, Wednesday

03:00-09:00; Corentin: Started with pointing on 1055+018, 2 reps on COSMOS project (little glitch of the tau for one scan) and finally another pointing. Moving to project nikaw-21 --> skydip [20220209s25] on 1418+546 because of the burst of opacity but then tau stayed stable ~0.23. Moving to project 199-16, pointing on 1418+546 and 6 reps on PSZ2G087 (18/24 --> 24/24 COMPLETED). Focus on 1633+382 and then 2 reps on PSZ2G084 (16/22 --> 18/22). [Retuning at 07:30: bad KIDs 36/5/41 for ar1/ar2/ar3]. Last pointing and moving on PSZ2G084 because it was the closest from the pointing source: 1 rep (18/24 --> 19/24) and I launched another rep but there was a failure with array 2 so the scan was canceled (need to update the README and put 19/24 instead of 20/24). [At 08:50 the arr2 crate went down suddenly; ./ crateColdReboot didn't work. Crate 2 restarted from PSE538: OK. DAQ restarted, retuning, bad KIDs 41/5/49 for ar1/ar2/ar3. Resuming observations].

09:00-15:00; Julien: Shift started by a technical issue with the array 2, followed by a power problem, all quickly solved (by 8:20 UTC). First project was nikaw-21 project toward NGC7027 (pointing, focus) to do calib_1scan on NGC7027, MWC349, and CRL2688 [scans 20220209s86, s87, and s88]. Good conditions, very stable (<2 Jy/beam peak-to-peak baseline variations and tau ~0.2). Moving to project 109-21. Several focus/pointing on 1757-240 to get a good beam, 1 repetition of B68 observation blocks, already getting low. Back to nikaw-21, doing poiting/focus/calib_1scan on Mars [s126]. Pointing/focus on 1749+096. Changing to project 109-21, pointing on 1923+210, then 2 repetitions on source L694-2, 1 pointing, 1 more repetition on L694-2. Opacity increases to ~0.4, so moving to project nikaw-21, trying to do pointing/focus on NGC7538. Repeated difficulties (too large and elongated beams), probably related to unrelaxed antenna which is likely to be heated by the Sun considering the clear weather.

15:00-21:00; Hélène: Rough start with distorted, but moderately large beams on 0316+413. The first 2 hours were devoted to multiple point-and-focus sequences away from the sun, with a skydip [scan s160] and a calib_1scan on three secondary calibrators (MWC349, NGC7027, and CRL2688 [scans s177, s178, and s179]), and on Uranus [scan s187]. Meanwhile, the antenna relaxed enough. Switched to program 122-16 (single source L1521e) for 4 scans at 60"/s and 2 scans at 45"/s. The beams were of better quality on the nearby pointing source (0439+360) than on Uranus, with FWHM ~ 17 to 18" at 2mm, and the focus was stable at -1.43 to -1.41 (according to the IDL RTA results). Signs of moderate anomalous refraction, acceptable sky rms (4 to 6 Jy/beam), tau of the order of 0.3 (maybe slightly too high for this program, the scan quality will need to be checked before charging the time). Relatively large pointing corrections, up to -8" in elevation. Then L1521e got too high, above 75 degrees. Switched to program 137-21 for 75 min on WB89-288. Uncertain focus (from -1.43, got to -0.64 and -1.03 in two successive focus sequences). Good stability (sky rms ~ 2.6 and 0.8 Jy/beam at 2mm in the first two scans), tau decreasing from ~0.3 to ~0.25.

21:00-03:00; Matthieu: After finishing the script of project 137-21 started during the previous shift, I performed calibrations on Uranus (pointing, focus, beam map, calib_1scan [scans s234, s235, and s236]). I then observed the NGC3184 of the IMEGIN project for 55min. I then switched to the COSMOS field and integrated (with small interruptions for pointing/focus) until the end of the shift. Overall, the weather was good with tau oscillating slowly between 0.25 and 0.2. It was a great night!

Feb 10th, Thursday

03:00-09:00; Emmanuel: [07:25: retuning, bad KIDs 44/5/45 for ar1/ar2/ar3; skydips: scans 20220210s25 and s69]

[Angel: At ~06:30 Emmanuel noticed that the beam size was about 17.3" x 19.1" at 2 mm and 12.1" x 18.9" (ar3; for unknown reasons, PIIC-QL monitor1 was deactivated sometime in the early morning) after a good/stable focus corrections. He tried with two additional sources with similar results. After notifying us, we checked for anomalies in the temperature of the back structure/quadrupod/ambient, among others, and we found no apparent reason to explain this effect. The atmosphere was nicely stable. A similar incident was reported by Emmanuel and Aishwarya on past Feb. 7th, same observing slot, with very low opacity and good stability. It's quite curious because this is precisely the time slot when the telescope is most relaxed. I took over from Emmanuel this morning and asked Salvador about the anomaly. He re-checked everything with the same conclusions. Following his suggestions, we inspected the main dish surface looking for dew depositions in the lower half that could induce beam distortions (according to the experience of Juan Peñalver). We found no evidence of such a thing. Focus and pointing on strong sources were checked again, even past sunrise. Pointing and focus corrections were quite stable, but the beam elongation remained: the best beam size got after that was about 17.5" x 18.3". We decided to continue with science projects. All your suggestions that could help to explain/fix this issue are welcome.]

09:00-15:00; Julien: The shift started after calib_1scan on NGC7027, MWC349 and CRL 2688 [scans 20220210s93, s94, and s95]. For Project 109-21 we observed L694-2 for 6 repetitions, with pointing and focus on 1923+210 every 1h. Good beam (17-18"), although a bit elongated. Good stability (<1 Jy/beam) but degrades when approaching noon. Switch to nikaw-21, focus/pointing on 0234+285, calib_1scan on Uranus [scan s128]. Beam was elongated, repeated focus and pointing. Switch to project 110-21, pointing on 0333+321 gives a rather circular but a bit large beam (18.34"x18.97" @2mm). Observing NGC1333-C7 for 6 repetitions until passing to the next observer, with pointing and focus on 0234+285 every hour. Stability continues degrading from ~3 Jy/beam when starting NGC1333-C7 until ~6 Jy/beam at the end of the shift.

15:00-21:00; Jean-Francois: Throughout the shift weather looked great (tau ~0.2-0.25 and stable according to piic rms). Programs 111-21 and 122-16 were observed. I began with pointing and focus on 3C84 (beam 17.6"x23.4") and then continue on 110-21 (source _C7) for 6 maps (66 minutes). Then pointing and focus on 3C84 (beam 17.5"x21.9") with the Sun opposite to the direction of observation (telescope back structure facing the Sun). Then switch to 122-16 (source L1521e above 45deg) for 4 scans at the mapping speed of 60"/sec, then 1 scan at 40"/sec and then 2 more scans but slowing at 30"/sec to pass successfully transit at 79.26 deg just below the elevation limit. I considered that the atmosphere was stable enough, apparently it was, for proper atmospheric fluctuation removal with this reduced speed during data reduction (something I'll be interested to check with Hélène). At beginning of the shift, beam was elongated at 17.6"x23.4" at 2mm but improved steadily with hours. It was 17.4"x18.7" one hour before sunset at 18h48 local and 17.5"x17.7" one hour after. Operations have been smooth thru VNCs for pako, IDL and piic. [Got a calib_1scan on CRL618 (s181)].

21:00-03:00; Nicolas: [Retuning at 21:05; bad KIDs 47/7/46 (ar1/ar2/ar3)]. Uranus, focus/pointing. [DAQ on ar2 crashes at 21:50 in the middle of a beam map. Crate 2 restarted from PSE538. OK. DAQ restarted, retuning: bad KIDs 44/5/48]. Relaunched the beam map after this, everything went smooth. Then calib_1scan on Uranus, skydip. Around 22h50 LT, switched to project 110-21: NGC1333-C7 is just nearby! Nice beams during the pointing on 0333+321. Did the last 3 needed rep on NGC1333-C7, this source is complete. After focus and pointing on 0736+017, now observing Cosmos (0h20 LT) until the end of the shift.

Feb 11th, Friday

03:00-09:00; Aishwarya: I began the shift with a pointing and focus on source 1055+018. The beam was circular at this point and of optimal size. I went on to perform 2 reps for the COSMOS field of project N2CLS. I changed to project 199-16 and performed a sequence of pointing and focus on source 1641+399 to get a good beam since at this point the beam had become elongated and also wider than optimal for the project. Two reps were performed on PSZ2G081 (199-16). At 6:10 the beam was still elongated but also much wider than optimal. A repeated sequence of focus and pointing was performed by changing sources which seemed to make the beam better but then high wind speeds caused observations to be halted till the end of the shift. [Stopped at ~07h: unbalanced motor currents due to the wind].

09:00-15:00; Angel: Telescope stopped because of high wind speed until 11:35. Resuming observations with NIKA2. Opacity is about 0.3 with fluctuations; the forecasts indicate that would increase. Retuning; bad KIDs 42/6/42 (ar1/ar2/ar3). Pointing and focus on 2251+153 for checking the beams. FWHMs of 18" x 18.4" (2 mm) and 13.1" x 13.4" (1 mm). Earlier elongation is practically absent. Beam broadening may be attributed to the exposition of the main dish to solar radiation while the telescope was in stow position. Opacity rises to 0.4. Regarding that and looking at the forecasts, we decided to switch to EMIR backup programs at 12:30. Starting with source 23219 of project 174-21 (2 iterations, ongoing).

15:00-21:00; Angelos (on-hold), Angel: EMIR backup program 174-21 continued: 3 iterations more on 23219. After an agreement with the B-FUN team, we changed to NIKA2-pol. DAQ restarted, HWP inserted and phased. Retuning: bad KIDs 46/6/52 (ar1/ar2/ar3).

21:00-23:45; B-FUN shift - Hamza, Quang, Philippe, Andrea: Good weather, but marginal opacity conditions with tau225 ~ 0.3-0.35. Reasonably stable atmosphere. We switched from EMIR to NIKA2-POL around 20:30 (LT) and took a beam map + calib1scan on Uranus (20220211s184, s185), then observe 1 sequence of 3 maps on OMC1 for B-FUN (20220211s194, s195, s196). This was followed by another calib1scan on Uranus (20220211s204) and a skydip (20220211s205). Due to the relatively high tau225~0.3 (too high for weaker B-FUN sources), we decided to switch back to EMIR after 23:30 (LT). [DAQ stopped at 23:52; switching to EMIR].

Feb 12th, Saturday

00:00-04:00; Angel: Restarting EMIR observations. Opacity dropping slowly from ~0.3. We got 4 loops on the targets 7926-high and 17370-high of project 170-21. Opacity level is about 0.27 and with reasonable chances of going down. Switching to NIKA2. DAQ started; retuning; bad KIDs 41/5/45 (ar1/ar2/ar3).

04:00-09:00; Frédéric: Switch to NIKA2 at 4:10 am., several pointing/focus scans. Start 160-16 at 5 am up to 7:20 am., done 4 repetitions on NGC3184. Readme file was updated. Switch to 137-21 at 07:23 pointing/focus, start obs on WB89-060. Observation launched at 08:10.

09:00-15:00; Angel: Continuing on WB89-060 (project 137-21); tau(225 GHz) about 0.25 and excellent stability: ~0.4 Jy/b (2 mm). Nice beam shape at 2 mm: FWHM 17.3" x 17.7". A set of calib_1scan maps were got on MWC349, CRL2688, and NGC7027 (scans s80, s81, and s82). A skydip after that (s83). Opacity is about 0.28 and tends to rise. Coming back to project 137-21 (this is the only project of the Pool that allows tau(225 GHz) < 0.5, stable atmosphere). Target WB89-031 was observed. The atmosphere is a bit unstable at times and the opacity still increasing: ~0.33 just after a skydip (s107). Switching to EMIR backup programs at 12:00. Project 174-21. Trying to get good pointing and focus without success. The atmosphere has become more unstable. Stopped at 13:15. Snowing since ~15h.

15:00-21:00; Angelos (on-hold). [Stopped. Snowing].

21:00-03:00; B-FUN team (on-hold). [It stopped snowing. Joaquín cleaned of snow/ice the subreflector and quadrupod at 22:30. Ice spots in selected zones of the main reflector. Deicing switched on. Monitoring antenna temperatures and main surface conditions].

Feb 13th, Sunday

03:00-09:00; Miguel, Angel (Marco & Antonio on-hold for NIKA2): Resuming observations. Opacity about 0.3 to 0.4. Starting EMIR backup project 170-21 (E090 and E230 bands). A bit unstable at the beginning. Targets observed: 55131-high (4 loops), 65716- high (8 loops), and 62323-high (2 loops); tau(225 GHz) ~0.3 at the end of the shift.

09:00-15:00; Juan, Angel: Continuing on project 170-21. Two more loops on 62323-high; tau(225 GHz) ~<0.3; evaluating possibilities for switching to NIKA2. Opacity would increase throughout the morning according to meteo models - confirmed by the actual opacity measurements. We decided to continue with EMIR. Unstable intervals. We completed 37963-high (4 loops) and started 41455-high (8 loops). Switching to project 174-21 (E090 band). Atmosphere has become more unstable. Opacity increased to 0.4-0.6. Observing target 23129 (1 repetition); tau triggered to ~1.6 in the last loop of this repetition. Telescope stopped at 14:24. Snowing slightly after that.

15:00-21:00; Angelos (on-hold). [Stopped. Snowing. Then, dense fog].

21:00-03:00; Hamza (on-hold). [Stopped. Dense fog].

Feb 14th, Monday

03:00-09:00; Laurence (on-hold for NIKA2). [Stopped. Dense fog. Pointing with EMIR tried by Joaquín at ~02:20. Stopping again. Snowing. Deicing switched on at 06:10].

09:00-15:00; Angel: Deicing switched off at 11:00. Telescope stopped until ~13h; tau<0.1 and a very stable atmosphere. Checking surfaces and NIKA2 setup. DAQ started, bad KIDs 28/4/47 (ar1/ar2/ar3). Updating inclinometers after the first pointing. Next pointing correction: 1.9" -6.3"; FWHM 17.3"x17.5" (2 mm) after a good focus. A bit enlarged at 1 mm: about 12". Getting calib_1scan maps on MWC349, NGC2027, and CRL2688 (scans 20220214s23, s24, and s25). Then a skydip (s26). Continuing on project 109-21: 2 reps. launched on L694-2. To be finished in the next shift.

15:00-20:00; Angelos: Opacity (tau) below 0.1 throughout all the shift. Continuing on project 109-21 until finishing started scans. Switching to project 110-21, 4 reps. (88 min) on B213-C16. After focus/pointing on CRL618, getting a @calib_1scan (s48), and then 1 rep. on L1524S (observed with a scan speed of 50"/sec due to high elevation). Switching to project 143-21 for 1 rep. on G202 until the end of the shift (20:00). [02:55: DAQ started for polarimetry; similar bad KIDs counts of morning retuning].

21:00-03:00; B-FUN shift - Philippe, Hamza, Andrea: Superb, stable weather conditions with tau225 ~ 0.08. Focus was stable at Fz = -0.50 +/- 0.1mm. We took two calib_1scans on Uranus (20220214s79 and s94), and observed 1 sequence of 3 maps in OMC1, 1 sequence of 3 maps on OMC2, 3 sequences of 2 maps on OMC3, and 2 sequences of 3maps + 2 extra maps on NGC2264 for the B-FUN project. Stopped observing with NIKA2-Pol at 02:45 LT after a calib_1scan on 0736+017 (20220215s10) to change over to NIKA2 TP. [02:55: DAQ started, bad KIDs 28/5/36 (ar1/ar2/ar3)].

Feb 15th, Tuesday

03:00-10:30; Juan. Very good weather, tau255 < 0.1 !!!, started with N2CLS: 2 reps in COSMOS after pointing and focus, switched to LPSZ: weird beams, repeated pointing and focus until got good beams, 8 reps in PSZ2G099.8, 1 rep PSZ2G091, 3 rep PSZ2G084. Ended with calibrations on secondary calibrators [s98, s99, s101, s103] and a skydip [s104].

DailyReportsNika2Pool080222 (last edited 2022-02-15 16:44:04 by NikaBolometer)